# An equivalence in random matrix and tensor models via a dually weighted intermediate field representation

### v1

@@ Appendix A The matrix $C_{2}$
for p=1,,Np=1,...,N. An explicit expression for $C_{m}$ in the large N limit was obtained in [abranches2025]. The matrix $C_{m}$, and particularly $C_{2}$, appears in different and seemingly unrelated contexts in matrix models and representation theory. Some examples are:
  • Character expansion. It is known [fulton1991representation] that for two given matrices $A$ and $B$ in ${\rm GL}(N)$, their traces and characters satisfy the following property:
@@
    where the first equality is obtained by expanding the l.h.s. in terms of the eigenvalues $a_{i}$ and $b_{i}$ of $A$ and $B$, respectively. The second equality comes from Cauchy’s identity (see equation A.13 in [fulton1991representation]). By choosing $B=C_{m}$, we get the following character expansion formula in the large-$N$ limit:
@@
    As shown in [difrancesco1992generatingfunctionfatgraphs], as well as an alternative proof in [abranches2025], this integral satisfies
@@
    In Section 4.5, we show a generalization of this. Namely, in Corolaries 4.3.3 and 4.3.4, we show that $\left\langle\chi_{r}(A)\right\rangle$ can be expressed in terms of $C_{1}$ in some non-Gaussian ensembles for any representation $r$ of ${\rm GL}(N)$.
@@ Appendix D Proofs
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,Q]]=(iN)^{-n}\sum_{m=0}^{\infty}\frac{i^{m}}{m!}\sum_{\mu\in S_{m}}{\rm Tr}_{[\sigma]}\left(\frac{\partial}{\partial B}\right){\rm Tr}_{[\mu]}(Q){\rm Tr}_{[\mu]}\bigl(PB\bigr)|_{B=0}.$ (D.17)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,Q]]=\frac{1}{n!}\,N^{-n}\sum_{\mu\in S_{n}}{\rm Tr}_{[\sigma]}\left(\frac{\partial}{\partial B}\right){\rm Tr}_{[\mu]}(Q){\rm Tr}_{[\mu]}(PB)\;.$ (D.18)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,Q]]=\frac{1}{n!}\,N^{-n}\sum_{\mu,\nu\in S_{n}}{\rm Tr}_{[\mu]}(Q){\rm Tr}_{[\sigma\nu^{-1}\mu^{-1}\nu]}(P)\;.$ (D.23)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,Q]]=N^{-n}\sum_{\mu\in S_{n}}{\rm Tr}_{[\mu]}(Q)\,{\rm Tr}_{[\mu\sigma]}(P),$ (D.24)
@@ Appendix E Equivalences for the quartic models
In this Appendix, we present an explicit application of the equivalences we proved for quartic complex matrix and tensor models. These examples are particularly instructive, as they show how the presence of the matrix $C_{2}$ (3.6) in the propagator effectively reduces a quartic interaction to a quadratic/Gaussian model.

### v2

@@ Acknowledgements
Conflicts of Interest
The authors have no conflict of interest to declare.
@@ Appendix A The matrix $C_{2}$
for p=1,,Np=1,...,N. An explicit expression for $C_{m}$ in the large N limit was obtained in [1]. The matrix $C_{m}$, and particularly $C_{2}$, appears in different and seemingly unrelated contexts in matrix models and representation theory. Some examples are:
  • Character expansion. It is known [22] that for two given matrices $A$ and $B$ in ${\rm GL}(N)$, their traces and characters satisfy the following property:
@@
    where the first equality is obtained by expanding the l.h.s. in terms of the eigenvalues $a_{i}$ and $b_{i}$ of $A$ and $B$, respectively. The second equality comes from Cauchy’s identity (see equation A.13 in [22]). By choosing $B=C_{m}$, we get the following character expansion formula in the large-$N$ limit:
@@
    As shown in [15], as well as an alternative proof in [1], this integral satisfies
@@
    In Section 5.5, we show a generalization of this. Namely, in Corolaries 5.3.3 and 5.3.4, we show that $\left\langle\chi_{r}(A)\right\rangle$ can be expressed in terms of $C_{1}$ in some non-Gaussian ensembles for any representation $r$ of ${\rm GL}(N)$.
@@ Appendix D Proofs
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,Q]]=(iN)^{-n}\sum_{m=0}^{\infty}\frac{i^{m}}{m!}\sum_{\mu\in S_{m}}{\rm Tr}_{[\sigma]}\left(\frac{\partial}{\partial B}\right){\rm Tr}_{[\mu]}(Q){\rm Tr}_{[\mu]}\bigl(PB\bigr)|_{B=0}.$ (D.17)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,Q]]=\frac{1}{n!}\,N^{-n}\sum_{\mu\in S_{n}}{\rm Tr}_{[\sigma]}\left(\frac{\partial}{\partial B}\right){\rm Tr}_{[\mu]}(Q){\rm Tr}_{[\mu]}(PB)\;.$ (D.18)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,Q]]=\frac{1}{n!}\,N^{-n}\sum_{\mu,\nu\in S_{n}}{\rm Tr}_{[\mu]}(Q){\rm Tr}_{[\sigma\nu^{-1}\mu^{-1}\nu]}(P)\;.$ (D.23)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,Q]]=N^{-n}\sum_{\mu\in S_{n}}{\rm Tr}_{[\mu]}(Q)\,{\rm Tr}_{[\mu\sigma]}(P),$ (D.24)
@@ Appendix E Equivalences for the quartic models
In this Appendix, we present explicit demonstrations of the equivalences we proved for quartic complex matrix (Theorem 5.3) and tensor models (Theorems 6.3 and 6.4). These examples are particularly instructive, as they show how the presence of the matrix $C_{2}$ (3.6) in the propagator effectively reduces a quartic interaction to a quadratic/Gaussian model.